OpenStudy (anonymous):

A farmer wants to enclose 135000 square yards of land in a rectangular plot. The material used for the front side costs $2 per yard, the material for the other three sides costs $1 per yard. Determine the dimensions that will minimize cost.

OpenStudy (campbell_st):

let l = length and w = width for area 135000 = lw or w = 135000/l (1) perimeter P = 2l + 2w substitute 1 from area formula P = 2l + 135000/l find dP/dl and thne let dP/dl = 0 and solve for l this will given the minimum length and subsequently minimum width by sustituting into (1)

OpenStudy (kropot72):

My calculation shows that the dimensions as follows will minimise cost: Front and rear lengths both equal 300 yards. Side lengths both equal 450 yards. The method follows.

OpenStudy (kropot72):

Let the front and rear lengths equal a yards each. Let the side lengths equal b yards each. Then a * b = 135000 and b = 135000/a Let cost = C C = 2a + a + 2b = 3a + 2b Substituting for b: C = 3a + 2*(135000/a) = 3a + 270000/a dC/dt = 3 - (270000/a^2) Put the differential coefficient = 0 to solve for a minimum: 3 - (270000/a^2) = 0 From which a^2 = 270000/3 = 90000 a = 300 yards by substitution in a * b = 135000 the value of b = 450 yards